# HESSO: Towards Automatic, Efficient, and User Friendly Any Neural Network Training and Pruning

### v1

@@ 4 Numerical Experiments
We numerically demonstrate the efficacy of HESSO across a wide range of applications, from low-level vision tasks such as super-resolution (Zhou et al., 2024), to high-level vision tasks object detection (Shi et al., 2020), as well as natural language processing tasks such as question answering (Rajpurkar et al., 2016) and the popular foundational large language models (Ding et al., 2023). The architectures used in these experiments encompass a variety of CNN benchmarks (Chen et al., 2023c) and transformers (Vaswani et al., 2017). These experiments involve training either from scratch or using a pre-trained checkpoint (when available) to validate the versatile application of HESSO-(CRIC).
In our experimental setup, we begin by directly applying HESSO due to its simplicity and resource efficiency, as it is capable of handling most applications and scenarios. If the direct application of HESSO encounters limitations, we then switch to HESSO-CRIC. It is important to note that HESSO-CRIC requires maintaining part of the model’s variables, resulting in slightly higher memory consumption. Across these experiments, HESSO-(CRIC) consistently achieves competitive, and in some cases, surpasses state-of-the-art results. Given its generality and ease of use, HESSO-(CRIC) stands out as one of the leading optimizers for joint training and pruning.
@@ 4.1 Recommended Experimental Setup
We recommend the following hyperparameter configurations for HESSO and HESSO-CRIC across varying applications and DNN architectures. For the target DNN to be trained and compressed, end-users likely already have a well-established training pipeline that enables the DNN to achieve high performance. To ensure ease of use, we suggest inheriting the hyperparameters in HESSO and HESSO-CRIC from the baseline training schema wherever there is overlap, such as in optimization variants and first- and second-order momentums.
This inheritance strategy should also be applied to other hyperparameters related to the training pipeline, such as training steps and learning rate schedules, though some slight adjustments may be needed for some applications. Specifically, adjustments may be needed due to the hybrid training process. We recommend beginning pruning at 1/10 of the total training steps and completing progressive pruning over another 1/10 of the total training steps. Because of the hybrid training stage, the learning rate schedule might require modification to ensure the DNN is sufficiently trained at a reasonably high learning rate after reaching the target group sparsity level.
Additionally, HESSO and HESSO-CRIC support training either from scratch or from a pre-trained checkpoint. For better performance and faster convergence, we recommend starting from a pre-trained status if such a checkpoint is available. We summarize the recommended hyperparameter selections and training strategies in Table 1. Remark that better hyperparameter setups or training strategies may exist for specific domain tasks to achieve superior performance.
For the remainder of the manuscript, we conduct the experiments according to the above recommenced criteria, unless otherwise specified. Additionally, all experiments were conducted on an NVIDIA A100 GPU with 80GB of memory.
@@
[…]
@@ 4.2 Super Resolution
We first selected the popular CARN architecture (Ahn et al., 2018) for the super-resolution task with a scaling factor of two, referred to as CARNx2. The benchmark DIV2K dataset (Agustsson and Timofte, 2017) was used for training, while Set14 (Zeyde et al., 2010), B100 (Martin et al., 2001), and Urban100 (Huang et al., 2015) datasets were employed for evaluation. Initially, we utilized OTO’s pruning dependency analysis to identify minimally removable structures and partitioned the trainable variables into pruning-zero-invariant groups. However, directly applying DHSPG or HESSO often led to significant performance degradation that was not reversible. This issue arose due to the architectural design, where the penultimate convolutional layer in CARN plays a crucial role in producing satisfactory visual results, making it a indispensable structure. Pruning minimally removable structures from this layer caused the remaining filters to fail in generating reasonable visual outcomes. The saliency score estimation in DHSPG and HESSO lacked the ability to avoid pruning such essential structures, resulting in irreversible performance collapse.
In OTOv2’s paper (Chen et al., 2023c), we manually excluded these indispensable structures from pruning. However, this manual identification is time-consuming and requires expert knowledge. To address this, we directly applied HESSO-CRIC to CARN and observed that it reliably identified these crucial structures as important groups, leading to a successfully pruned model. Specifically, following (Agustsson and Timofte, 2017), we used Adam as the optimization variant in DHSPG, HESSO, and HESSO-CRIC, leveraging both first- and second-order momentum to estimate the gradient. As shown in Table 2, when manually excluding indispensable structures, both DHSPG and HESSO significantly reduced FLOPs and parameters by approximately 33% to 80%, with negligible PSNR degradation. HESSO-CRIC achieved a better trade-off between FLOP reduction and PSNR, as demonstrated by exhibiting the frontier curve under varying pruning ratios. Furthermore, it can automatically identify indispensable structures, preventing irreversible performance collapse.
@@
HESSO-CRIC Automatic 40% 40.4% 40.1% 33.3 32.0 31.4
@@
(a) Low resoluted image.
(b) Group Sparsity = 20%.
(c) Group sparsity = 30%.
(d) High resoluted image.
(e) Group Sparsity = 50%.
(f) Group Sparsity = 60%.
@@ 4.3 Question and Answering
HSPG (Chen et al., 2021) 91.0% 84.1%
HSPG (Chen et al., 2021) 66.7% 82.0%
@@
Later, we compare HESSO-(CRIC) with DHSPG, HSPG, and a representative proximal method (ProxSSI) for pruning a transformer model Bert(Vaswani et al., 2017), evaluated on the SQuAD question-answering benchmark(Rajpurkar et al., 2016). It is important to note that proximal methods have been standard algorithms for solving sparse optimization problems for decades. However, recent research has shown that they are not effective at exploring sparsity while maintaining model performance in deep learning applications (Dai et al., 2023).
Following previous work (Chen et al., 2023c), we excluded the embedding layer from pruning. As shown in Table 3, HESSO, HESSO-CRIC, and DHSPG perform competitively on this task in terms of parameter reduction while maintaining F1 scores. However, DHSPG only achieves these results after extensive hyper-parameter tuning, which is neither efficient nor convenient. While HSPG performs comparably, its tendency to penalize all variables toward zero severely restricts the optimization search space, leading to suboptimal performance. Proximal methods, contrarily are not comparable with the HESSO and HSPG families.
4.4 Object Detection
Next, we tested HESSO on the popular YOLO (Redmon et al., 2016) object detection model using the COCO benchmark dataset (Lin et al., 2014). Table 4 presents the structural pruning results for YOLOv5l (Jocher et al., 2022) using HESSO and HESSO-CRIC. It is important to note that we selected YOLOv5l to facilitate comparisons with other existing benchmark methods. Starting from a pretrained YOLOv5l model, we applied HESSO and HESSO-CRIC with a target group sparsity of 30%, resulting in a sub-network containing 49% of the original parameters. This allows for direct comparison with benchmarks that retain 50% of the model’s parameters. The results show that a single run of HESSO and HESSO-CRIC achieved significantly higher Mean Average Precision (mAP) compared to other pruning approaches, which often require more complex, multi-stage procedures.
Table 4: Structurally pruning Yolov5l on COCO.
  Method # of Params $\text{mAP}_{0.5}$ $\text{mAP}_{0.5:0.95}$
Baseline 100% 66.31% 47.71%
[…]
TCFP (Jeon et al., 2022) 50% 61.8% 42.7%
HESSO (30% group sparsity) 49% 63.1% 44.4%
HESSO-CRIC (30% group sparsity) 49% 63.1% 44.5%
(a) Pretrained YOLOv5l.
(b) Group Sparsity = 30%.
(c) Pretrained YOLOv5l.
(d) Group Sparsity = 30%.
Figure 5: Visual examples of pruned YOLOv5l.
@@ 4.5 Large Language Model
Finally, we evaluated HESSO-(CRIC) on large language models (LLMs). Since both HESSO and HESSO-CRIC utilize full gradient information, we focused on LLMs with fewer than 3 billion parameters, such as the representative Phi-2-2.7B(Microsoft, 2023), to ensure that a single 80GB GPU is sufficient, without requiring tensor parallelism (Ding et al., 2023). Our experimental setup followed that of LoRAShear (Chen et al., 2023a).
@@
In contrast, HESSO-CRIC was able to automatically bypass these crucial structures, enabling effective and successful pruning. We then compared its results with SliceGPT (Ashkboos et al., 2024) across several popular benchmarks using the LM-Evaluation-Harness (Gao et al., 2021). Our findings indicate that HESSO-CRIC consistently outperforms SliceGPT at varying pruning ratios, with performance improvements becoming more pronounced as the pruning ratio increases. We have the perspective that the hybrid training scheme in HESSO serves a similar purpose by achieving computational invariance between pruned sub-network and full DNN, maintaining the same objective function value during progressive pruning. Ours is architecture agnositic and easy to apply onto various applications. However, SliceGPT requires significant manual efforts to introduce addition auxiliary structures and restricts onto transformer architecture. Furthermore, HESSO-CRIC offers the added advantage of a more reliable identification mechanism, excluding indispensable structures and better preserving performance, particularly at higher pruning ratios.
@@ 5 Conclusion
In this work, we have introduced HESSO, a novel Hybrid Efficient Structured Sparse Optimizer, designed to address the challenges of structurally pruning deep neural networks (DNNs) while maintaining their performance. HESSO streamlines the pruning process by incorporating a progressive pruning strategy guided by saliency scores, a hybrid training approach that transfers knowledge from redundant structures to essential ones, and a user-friendly interface suitable for a wide range of training applications. Additionally, we proposed the Corrective Redundant Identification Cycle (CRIC), a mechanism that mitigates the risk of irreversible pruning-induced performance collapse by reliably identifying and preserving indispensable structures within the network.
Our extensive numerical experiments across various domains, including computer vision and natural language processing, demonstrate that HESSO achieves competitive performance to state-of-the-art benchmarks, adapting to diverse architectures. Furthermore, the integration of CRIC into HESSO not only prevents performance degradation but also enhances the optimizer’s capability to deliver superior results in scenarios where traditional pruning methods falter.
@@
Appendix A Appendix
Proof of Theorem 3.2
@@ Appendix A Appendix
The statement is equivalent to that the violating cycle line 6-19 in Algorithm 2 terminates within finite number of steps. For convenience, we denote $\mathcal{V}_{l}$ as the violating set at $l$th cycle. The statement then becomes that there exists an $L<\infty$, such that $\mathcal{V}_{L}=\emptyset$. We now prove it as a two-step fashion.

### v2

@@ 4 Numerical Experiments
We numerically demonstrate the efficacy of HESSO across a wide range of applications, from high-level vision tasks including image classification (He et al., 2016), to low-level vision tasks such as super-resolution (Zhou et al., 2024), and object detection (Shi et al., 2020), as well as natural language processing tasks such as question answering (Rajpurkar et al., 2016) and the popular foundational large language models (Ding et al., 2023). The architectures used in these experiments encompass a variety of CNN benchmarks (Chen et al., 2023c) and transformers (Vaswani et al., 2017). These experiments involve training either from scratch or using a pre-trained checkpoint (when available) to validate the versatility of HESSO-(CRIC). Furthermore, we provided ablation studies of CRIC over different saliency scores in Section 4.7, hyper-parameter tuning effort studies in Section 4.8, and computational complexity analysis in Appendix C.
@@ 4.1 Recommended Experimental Setup
We recommend the following hyperparameter configuration in Table 1 for HESSO and HESSO-CRIC across various applications and DNN architectures. For the target DNN to be trained and compressed, end-users likely already have a well-established training pipeline that allows the DNN to achieve high performance. To enhance usability, we recommend inheriting the hyperparameters in HESSO and HESSO-CRIC from the baseline training schema wherever there is overlap, such as with optimizer variants and first- and second-order momentum.
@@
Optimizer variant HESSO-(CRIC) Inherit as the baseline optimizer. Currently support {SGD, Adam, AdamW}.
Group sparsity HESSO-(CRIC) Set upon the target pruned model size. If all variables could be pruned, the pruned model size could be approximately equal as quadratic of the density level. In addition, a randomly pruned model could be obtained by OTO’s APIs.
First-order momentum HESSO-(CRIC) Inherit as the baseline optimizer’s first-order momentum.
Second-order momentum HESSO-(CRIC) Inherit as the baseline optimizer’s second-order momentum.
Weight-decay HESSO-(CRIC) Inherit as the baseline optimizer’s weight-decay.
Initial learning rate HESSO-(CRIC) Inherit as the baseline optimizer’s initial learning rate.
Saliency Score Criteria HESSO-(CRIC) By default equally considering the scores in Section 3.1.
Start pruning step HESSO-(CRIC) Set up as 1/10 of total training steps.
Pruning steps HESSO-(CRIC) Set up as 1/10 of total training steps.
Pruning periods HESSO Empirically suggest to set as 10.
Sampling steps HESSO-CRIC Empirically suggest to set as 10.
Learning rate scheduler Training Inherit as the baseline training, yet might need adjustments in some application to ensure the model after reaching target group sparsity is sufficiently trained under relatively large learning rate.
Total training steps Training Inherit as the baseline training and adjust upon the learning rate scheduler.
Start training from scratch or pretraining checkpoint Training Both are supported. For better performance, recommend to start from pretraining checkpoint if available.
This inheritance strategy should also be applied to other hyperparameters related to the training pipeline, such as training steps and learning rate schedules, though some slight adjustments may be needed for some applications due to the hybrid training process. We recommend beginning pruning at 1/10 of the total training steps and completing progressive pruning over another 1/10 of the total training steps. Because of the hybrid training stage, the learning rate schedule might require modification to ensure the DNN is sufficiently trained at a reasonably high learning rate after reaching the target group sparsity level.
Additionally, HESSO and HESSO-CRIC support training either from scratch or from a pre-trained checkpoint. For better performance and faster convergence, we recommend starting from a pre-trained status if such a checkpoint is available. We summarize the recommended hyperparameter selections and training strategies in Table 1. Remark that better hyperparameter setups or training strategies may exist for specific domain tasks to achieve superior performance. For the remainder of the manuscript, we conduct experiments according to the above recommenced criteria.
@@ 4.2 Super Resolution
HESSO-CRIC Automatic 40% 40.4% 40.1% 33.3 32.0 31.5
@@
We then selected the popular CARN architecture (Ahn et al., 2018) for the super-resolution task with a scaling factor of two, referred to as CARNx2. The benchmark DIV2K dataset (Agustsson and Timofte, 2017) was used for training, while Set14 (Zeyde et al., 2010), B100 (Martin et al., 2001), and Urban100 (Huang et al., 2015) datasets were employed for evaluation. Initially, we utilized OTO’s pruning dependency analysis to identify minimally removable structures and partitioned the trainable variables into pruning-zero-invariant groups. However, directly applying DHSPG or HESSO led to significant performance degradation that was not reversible. This issue stems from the architectural design, where the penultimate convolutional layer is critical for generating satisfactory visual results, making it an indispensable structure. Pruning this layer caused the remaining filters to fail in generating reasonable visual outcomes. However, the saliency score deems them as redundant due to significant approximation errors and thus, results in irreversible performance collapse.
OTOv2 (Chen et al., 2023c) manually excluded these indispensable structures from pruning. However, this manual identification is time-consuming and requires expert knowledge. To address this, we applied HESSO-CRIC to CARN and observed that it automatically identified these crucial structures as important groups, leading to a successfully high-performing pruned model. As shown in Table 2, when manually excluding indispensable structures, both DHSPG and HESSO significantly reduced FLOPs and parameters by approximately 33% to 80%, with negligible PSNR degradation. HESSO-CRIC achieved a better trade-off between FLOP reduction and PSNR, as demonstrated by exhibiting the frontier curve under varying pruning ratios. Visual examples shown in Figure 6 at Appendix D further cross-verify the performance preservation by our approaches.
4.3 Image Classification
Figure 4: ResNet50 on ImageNet.
We first conducted the benchmark ResNet50 (He et al., 2016) on ImageNet. As displayed in Figure 4, HESSO-CRIC roughly exhibits a Pareto frontier in terms of top-1 accuracy and FLOPs reduction under various group sparsities from 40% to 70%. HESSO and DHSPG perform competitively in this application. Moreover, all of then could produce structurally pruned sub-networks associated with smaller size fewer FLOPs, and higher accuracy compared to most of the existing approaches (Huang and Wang, 2018, Zhou et al., 2019, Ding et al., 2021a, Yang et al., 2019, You et al., 2019, Zhou et al., 2019). These results well validate the efficacy of the newly proposed joint pruning and training optimizer on this popular structured pruning benchmark.
We further employ HESSO-(CRIC) to structurally prune a pretrained OFA network (Cai et al., 2020) on the benchmark ImageNet (Deng et al., 2009). The OFA network was produced by searching from a MobileNetV3 based super-network and could achieve 80.0% top-1 test accuracy on ImageNet. We find that both HESSO-(CRIC) could effectively discover pruned sub-networks with similar size and MACs while with higher performance than other OFA networks, i.e., 78.6% and 78.2% versus 76.9% test accuracy.
Table 3: Structurally pruning MobileNet Search Space.
  Method # of Params (M) MACs (M) Top Acc-1 (%)
[…]
MobileNetV3-Large (Howard et al., 2019) 5.4 219 75.2
OFA # 75 (Cai et al., 2020) 5.81 230 76.9
[…]
HESSO-CRIC 5.71 225 78.6
4.4 Object Detection
Table 4: Structurally pruning Yolov5l on COCO.
  Method # of Params $\text{mAP}_{0.5}$ $\text{mAP}_{0.5:0.95}$
  Baseline 100% 66.31% 47.71%
[…]
TCFP (Jeon et al., 2022) 50% 61.8% 42.7%
HESSO (30% group sparsity) 49% 63.1% 44.4%
HESSO-CRIC (30% group sparsity) 49% 63.1% 44.5%
Next, we tested HESSO on the popular YOLO (Redmon et al., 2016) object detection model using the COCO benchmark dataset (Lin et al., 2014). Table 4 presents the structural pruning results for YOLOv5l (Jocher et al., 2022). Note that we selected YOLOv5l to facilitate comparisons with other existing benchmarks. We applied HESSO and HESSO-CRIC with a target group sparsity of 30%, resulting in a sub-network containing 49% of the original parameters. This allows for direct comparison with benchmarks that retain 50% of the model’s parameters. The results show that a single run of HESSO and HESSO-CRIC achieved significantly higher Mean Average Precision (mAP) compared to other pruning approaches, which often require more complex, multi-stage procedures. For further visualization, additional details can be found in Figure 6 in Appendix D.
@@ 4.5 Question and Answering
Later, we compare HESSO-(CRIC) with DHSPG, HSPG, and a representative proximal method ProxSSI (Deleu and Bengio, 2021) for pruning a transformer model Bert (Vaswani et al., 2017), evaluated on the SQuAD question-answering benchmark (Rajpurkar et al., 2016). It is important to note that proximal methods have been standard algorithms for solving sparse optimization problems for decades. However, they are not effective at exploring sparsity while maintaining model performance in deep learning applications (Dai et al., 2023).
As shown in Table 5, HESSO, HESSO-CRIC, and DHSPG perform competitively on this task in terms of parameter reduction while maintaining F1 scores. However, DHSPG achieves these results after extensive hyper-parameter tuning, which is not convenient. HSPG penalizes all variables toward zero which severely restricts the optimization search space, leading to suboptimal performance. ProxSSI additionally lacks sufficient sparsity exploration capacity, being not comparable.
@@
HSPG (Chen et al., 2021b) 91.0% 84.1%
HSPG (Chen et al., 2021b) 66.7% 82.0%
@@ 4.6 Large Language Model
Finally, we evaluated HESSO-(CRIC) on large language models (LLMs). Since both HESSO and HESSO-CRIC utilize full gradient information, we focused on LLMs with fewer than 3 billion parameters, such as the representative Phi-2-2.7B (Javaheripi et al., 2023), to ensure that a single 80GB GPU is sufficient, without requiring tensor parallelism (Ding et al., 2023). Our experimental setup followed that of LoRAShear (Chen et al., 2023a).
@@
HESSO-CRIC was able to automatically bypass these crucial structures, enabling effective and successful pruning. We then compared with SliceGPT (Ashkboos et al., 2024), LLM-Pruner (Ma et al., 2023), LoraShear (Chen et al., 2023a) and LoraPrune (Zhang et al., 2023) across several popular benchmarks. Our findings indicate that HESSO-CRIC consistently outperforms them at varying pruning ratios, with performance improvements becoming more pronounced as the pruning ratio increases. This is because LLM-Pruner, LoRA-Prune, and LoRAShear are LoRA-based techniques. Lora primarily focuses on fine-tuning well-trained models and is less effective in capturing knowledge for underfitted models, such as pruned LLMs.
@@
4.7 Ablation Studies of CRIC on Saliency Scores
The default format of CRIC primarily targets the most commonly used saliency scores that are sensitive to approximation errors caused by distances to the origin. For saliency scores with such higher sensitivities, CRIC’s multiple sampling strategy—gathering information along the direction toward the origin—and its voting mechanism over historical statistics can effectively mitigate these identification issues.
To validate this, we have included a new ablation study for CRIC to demonstrate its improvements across varying saliency scores. As shown in the results, for commonly used saliency scores, CRIC effectively improves performance. However, magnitude and average magnitude benefits less from CRIC due to the persistence of large approximation errors, even as the groups of iterates move closer to the origin.
Table 7: Ablation Studies of CRIC on Zero-Shot Pruning Phi2.
Magnitude Avg Magnitude Cosine Similarity 1st Taylor 2nd Taylor
  No CRIC CRIC No CRIC CRIC No CRIC CRIC No CRIC CRIC No CRIC CRIC
  Perplexity↓ 629.1 489.4 713.5 644.6 525.5 53.4 438.3 28.6 378.2 37.1
Furthermore, for saliency scores whose approximation errors are not dependent on the distance to the origin, the philosophy of CRIC can still be applied with proper adaptations. In such cases, it is critical to analyze the root causes of the approximation errors for the given saliency scores. Based on these root causes, CRIC’s multiple sampling strategy can be adjusted to collect more targeted signals, thereby reducing identification errors in these scenarios.
4.8 Comparative analysis of hyper-parameter tuning efforts
The key advantage of HESSO-(CRIC) over HSPGs in the OTO series lies in its white-box optimization design. Unlike HSPGs, which are black-box optimizers requiring extensive task-specific hyper-parameter tuning for optimal performance, HESSO-(CRIC) significantly reduces this sensitivity by design. To highlight this difference, we present a comparative analysis of the total number of training recipes required for three shared applications:
Table 8: Sparse optimization related hyper-parameter recipe comparisons.
HESSO-(CRIC) DHSPG
  Super-Resolution CARNx2 General Recipe as described in Table 5 of manuscript. Recipe #1: $\lambda=10^{-2}$, $\lambda_{amplify}=20$, $\epsilon=0.0$, etc.
  Image-Classification ResNet General Recipe as described in Table 5 of manuscript. Recipe #2: $\lambda=10^{-3}$, $\lambda_{amplify}=2$, $\epsilon=0.95$, etc.
  Question-Answering Bert General Recipe as described in Table 5 of manuscript. Recipe #3: $\lambda=10^{-3}$, $\lambda_{amplify}=2$, $\epsilon=0.0$, etc.
  Total # of training recipes 1 3
As shown in Table 8, HESSO-(CRIC) achieves competitive or superior performance using a single general-purpose recipe, whereas DHSPG requires distinct task-specific hyper-parameter settings for each application.
Additionally, this comparison focuses only on hyper-parameters specific to sparse optimizers. Black-box optimizers like HSPGs inherently manage sparsity exploration processes, which demand further tuning of broader training parameters, such as learning rate schedules and the number of epochs. In contrast, the white-box design of HESSO-(CRIC) avoids such complexities, offering a more user-friendly, efficient, and practical solution.
@@ 5 Conclusion
In this work, we introduced HESSO-(CRIC), a novel Hybrid Efficient Structured Sparse Optimizer tailored for pruning deep neural networks while preserving performance. By combining a hybrid training strategy with explicit, progressive pruning control, and the Corrective Redundant Identification Cycle (CRIC), HESSO-(CRIC) effectively tackles challenges such as tuning efforts, user difficulty, and irreversible performance degradation. Our experiments across diverse domains show that it not only competes with but often surpasses state-of-the-art methods.
@@
Appendix A Proof of Theorem 3.3
@@
The statement is equivalent to that the violating cycle line 5-17 in Algorithm 2 terminates within finite number of steps. For convenience, we denote $\mathcal{V}_{l}$ as the violating set at $l$th cycle. The statement then becomes that there exists an $L<\infty$, such that $\mathcal{V}_{L}=\emptyset$. We now prove it as a two-step fashion.
@@
Appendix B Saliency Score
$\begin{split}\text{score}_{\text{mag}}([\bm{x}]_{g})&\leftarrow\left\|[\bm{x}]_{g}\right\|_{2},\\ \text{score}_{\text{mag}}([\bm{x}]_{g})&\leftarrow\frac{\text{score}_{\text{mag}}([\bm{x}]_{g})}{\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{mag}}([\bm{x}]_{g})}.\end{split}$ (4)
$\text{score}_{\text{avg-mag}}([\bm{x}]_{g})$ $\leftarrow\frac{\left\|[\bm{x}]_{g}\right\|_{2}}{|\sqrt{|g|}|},$ (5)
$\text{score}_{\text{avg-mag}}([\bm{x}]_{g})$ $\leftarrow\frac{\text{score}_{\text{avg-mag}}([\bm{x}]_{g})}{\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{avg-mag}}([\bm{x}]_{g})}.$
$\begin{split}\text{score}_{\text{cosine}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\frac{[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}}{(\left\|[\bm{x}]_{g}\right\|\left\|[\nabla\!f(\bm{x})]_{g}\right\|)},\\ \text{score}_{\text{cosine}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\frac{\text{score}_{\text{cosine}}([\bm{x}]_{g})}{\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{cosine}}([\bm{x}]_{g})}.\end{split}$ (6)
Taylor Importance. To further quantitatively approximate the effect of projecting the parameter group $[\bm{x}]_{g}$ onto zero on the objective function, we can employ the Taylor expansion. Taylor expansion could estimate the impact of small changes in the parameters on the function value, allowing us to consider varying orders of Taylor importance. In particular, the first-order Taylor expansion provides a linear approximation of the objective function around the current parameter point. The impact of setting $[\bm{x}]_{g}\to\bm{0}$ can be estimated by the dot product of the gradient and the change in parameters. It helps identify groups whose removal likely decrease objective function.
$\begin{split}\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow|f(\bm{x})-f(\bm{x}|[\bm{x}]_{g}\to\bm{0})|\approx|[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}|,\\ \text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\frac{\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})}{\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})}.\end{split}$ (7)
$\begin{split}\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow|f(\bm{x})-f(\bm{x}|[\bm{x}]_{g}\to\bm{0})|\approx[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}+\frac{1}{2}[\bm{x}]_{g}^{\top}[\nabla\!^{2}f(\bm{x})]_{g}[\bm{x}]_{g},\\ \text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\frac{\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})}{\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})}.\end{split}$ (8)
Appendix C Computational Cost Analysis
In this section, we present the time and space complexities of HESSO-(CRIC). For ease of presentation, we introduce several notations in Table 9.
Table 9: Notations.
Symbol Definition Remark
  $N$ of trainable variables with gradient
  $\mathcal{G}$ The set of parameter groups The common setup could be pruning/erasing zero-invariant groups.
  $|\mathcal{G}|$ The size of $G$ Typically negligible compared to $N$, see the below table.
  $T$ of training steps
  $T_{ht}$ of hybrid training steps Set as $T_{ht}=T/10$ in our generic recipe.
  $P$ of pruning periods Set as $P=10$ in our generic recipe.
  $S$ of sampling steps in CRIC Set as $S=10$ in our generic recipe.
  $C$ of cycles in CRIC Empirically terminates within 10 cycles.
HESSO-(CRIC) requires additional time and space complexities while the additions are negligible. In our numerous realistic applications besides the presented academic benchmarks, HESSO-(CRIC) are quite efficient, typically as efficient as standard training via vanilla optimizers. Detailed complexity results are presented in Table 10.
Table 10: Space and Time Complexity Comparison.
Optimizer Variant Space Complexity (Peak) Time Complexity Space Complexity Projected onto Phi2 Time Complexity Projected onto Phi2
  SGD Standard $O(2N)$ $O(NT)$ $O(2N)$ $O(NT)$
  HESSO SGD $O(2N+\|G\|)$ $O(NT+\|G\|P)$ $O(2.00015N)$ $O(NT+1.5\times 10^{-3}N)$
  HESSO-CRIC SGD $O(2N+\|G\|S)$ $O(NT+\|G\|P+\|G\|SC)$ $O(2.0015N)$ $O(NT+1.515\times 10^{-1}N)$
  Adam/AdamW Standard $O(3N)$ $O(2NT)$
  HESSO Adam/AdamW $O(3N+\|G\|)$ $O(2NT+\|G\|P)$ $O(3.00015N)$ $O(2NT+1.5\times 10^{-3}N)$
  HESSO-CRIC Adam/AdamW $O(3N+\|G\|S)$ $O(2NT+\|G\|P+\|G\|SC)$ $O(3.0015N)$ $O(2NT+1.515\times 10^{-1}N)$
Appendix D Supplementary Pictures
(a) Pretrained YOLOv5l.
(b) Group Sparsity = 30%.
(c) Pretrained YOLOv5l.
(d) Group Sparsity = 30%.
Figure 5: Visual examples of pruned YOLOv5l.
(a) Low resoluted image.
(b) Group Sparsity = 20%.
(c) Group sparsity = 30%.
(d) High resoluted image.
(e) Group Sparsity = 50%.
(f) Group Sparsity = 60%.